Can someone help me with this problem?
I need to evaluate the following expression if a = -2, b = 5, c = -4, and d=6
d^2 - b^2
Thanks,
Sher

Answer by checkley71(8403)

d^2=6*6 & b^2=5*5 thus (6*6)-(5*5) or 36-25=9